# AI-assisted Programming May Decrease the Productivity of Experienced Developers by Increasing Maintenance Burden

### v2

@@ 3.1 Research Design
Figure 5: The timeline of our study period: the technical preview of GitHub Copilot was launched on June 29, 2021, initially endorsing five programming languages – Python, JavaScript, Ruby, TypeScript, and Go. We designed our study to include the 12 months before and the 12 months after Copilot’s introduction.
@@
In June 29, 2021, developed on OpenAI’s GPT-3 model, GitHub launched a technical preview version of Copilot – the very first AI pair programmer. This early version of GitHub Copilot was not open to the public and endorsed only five programming languages: Python, JavaScript, Ruby, TypeScript, and Go. [^18] https://github.blog/news-insights/product-news/introducing-github-Copilot-ai-pair-programmer/ Copilot was later launched to the public in June 2022, with contributors required to pay a monthly fee to subscribe to the AI pair programming service. With the public release in June 2022, Copilot gradually added endorsement for more languages such as, C and Java. As shown in Figure 5, we define our observation period as 12 months before and 12 months after its introduction. The measures were aggregated to create a monthly panel dataset spanning from July 2020 to July 2022.
@@
Our study leverages this natural experiment created by the launch of GitHub Copilot, an AI-powered LLM designed to assist with coding. Specifically, we take advantage of the early stage Copilot’s limited language endorsement, which included languages of Python, JavaScript, Ruby, TypeScript, and Go, while excluding others comparable languages of R, C, C#, C++, Java, PHP and Scala. We select these five language as the non-Copilot-endorsed languages because of their comparable functionality and also they are the most frequently used languages for Microsoft-owned repositories. Our identification strategy by programming language between treatment and control group are similar to those in past studies (Yeverechyahu et al. 2024).
@@
Anecdotal evidence [^19] https://www.mckinsey.com/industries/technology-media-and-telecommunications/our-insights/thomas-dohmke-on-improving-engineering-experience-using-generative-ai and our interviews of Microsoft employees indicates that during the technical preview, access to Copilot was restricted to selected GitHub contributors, specifically employees of Microsoft/GitHub and maintainers of popular repositories. This restricted access suggests that the primary contributors of Copilot during the technical preview were likely Microsoft and GitHub employees, ensuring that the tool was stress tested (Dog Fooding; is Microsoft speak for internal use of their own software [^20] https://devblogs.microsoft.com/oldnewthing/20110802-00/?p=10003) internally before its broader release.
@@
Since Copilot usage cannot be identified at the individual level, we focus on Microsoft-owned repositories and contributors who actively contributed to these repositories during the observation period. By doing so, we aim to capture changes in OSS contributor behaviour driven by Copilot, as contributors to Microsoft-owned repositories are more likely to have free access to the tool.
@@
We use GitHub’s API service to collect all the data for this research, enabling efficient and precise querying of repository/individual activities. This approach enables us to gather detailed information on measurements such as pull requests, commits, reviews, authors, and repository/individual metadata, providing a comprehensive dataset for analyzing the impact of AI-assisted coding on OSS project development and maintenance.
We report the summary statistics for key variables used in our analysis. For the project level analysis, the binary variable Copilot indicates the treatment (treated repository*post copilot), with a mean of 0.30 and a standard deviation of 0.46. The mean log of lines of code added is 1.33 (SD = 2.88). The mean log of the commit count is 0.47 (SD = 1.05) and the mean log of PRs is 0.36 (SD = 0.83). The average log of PR rework – defined as the number of commits added to a PR after its submission – is 0.27 with a standard deviation of 0.89.
For the individual level analysis, the binary indicator for Copilot indicates the treatment (treated contributor*post Copilot), with a mean of 0.366 and a standard deviation of 0.482. The log-transformed number of commits submitted by each contributor per month has a mean of 1.75 (SD = 1.69), ranging from 0 to a maximum of 10.12. The number of PRs submitted, also log-transformed, has a mean of 0.96 (SD = 1.16), with a maximum value of 6.20. The log number of PR reviews conducted has a mean of 0.79 (SD = 1.21), reaching a maximum of 6.03. Lastly, the number of repositories where users conducted PR reviews, log-transformed, has a mean of 0.45 (SD = 0.66), with a maximum value of 5.21.
@@
3.2 Statistical analysis
To estimate the effect of Copilot on repositories and individuals, we used two difference in difference (DiD) regression model. The project level DiD model (repo i month t) is provided below:
@@ 3.2 Statistical analysis
where $y_{i,t}$ refers to the outcome measures (development and maintenance) for project / individual $i$ in month $t$. $Copilot_{i,t}$ is the independent variable and a binary indicator that turns to 1 when the Copilot is released and functioned as our treatment. $CoreContributor_{i}$ is the moderator and a binary indicator that turns to 1 when one individual is identified as core contributor by pretreatment code contribution behaviour. The project-level / individual-level fixed effects are represented by $\gamma_{i}$, and $\delta_{t}$ represents the monthly fixed effects. $\epsilon_{i,t}$ is the robust standard error clustered at the project / individual level to account for the potential heteroskedasticity of the errors.
@@ 3.3 Matching results
As a robustness check, we employed Coarsened Exact Matching (CEM) to mitigate concerns about potential selection bias and ensure a more balanced comparison between treatment and control groups. CEM allows us to pre-process the data by matching units on a set of covariates that may influence both treatment assignment and outcomes. We matched repositories based on pretreatment code development characteristics, including Code Added, Commits, and Pull Request. After matching, the repository level dataset consisting of n = 2,510 repositories, with 1,486 in the treatment group and 1,024 in the control group. The treatment and control groups were more closely aligned in their baseline characteristics, reducing imbalance across key variables. We then re-estimated the DiD models using the matched sample. The results are listed in the table below. The results remain consistent with our main findings, reinforcing the conclusion that the integration of GitHub Copilot is associated with increased rework and maintenance related activities.
@@
Concept: Development Maintenance

### v3

@@ 3.1 Research Design
Figure 2: The timeline of our study period
@@
In June 29, 2021, developed on OpenAI’s GPT-3 model, GitHub launched a technical preview version of Copilot [^10] In September 2020, Microsoft gained exclusive access to OpenAI’s GPT3. This access allowed Microsoft to repurpose and modify the model for code generation, leading to the development of GitHub Copilot. In June 2021, GitHub Copilot was launched as a technical preview. Anecdotal evidence and our interviews of Microsoft employees indicates that during the technical preview, access to Copilot was restricted to selected GitHub users, specifically employees of Microsoft/GitHub and maintainers of popular projects. This restricted access suggests that the primary users of Copilot during the technical preview were likely Microsoft and GitHub employees, along with selected Github uses who volunteered for the beta testing the software., an AI-powered LLM designed to assist with coding. This early version of GitHub Copilot was not open to the public and endorsed five programming languages: Python, JavaScript, Ruby, TypeScript, and Go. [^11] https://github.blog/news-insights/product-news/introducing-github-Copilot-ai-pair-programmer/ Copilot was later launched to the public in June 2022, with contributors required to pay a monthly fee to subscribe to the AI pair programming service. With the public release in June 2022, Copilot gradually added endorsement for more languages such as, C and Java. As shown in Figure 2, we define our observation period as 12 months before and 12 months after its introduction as technical preview. The measures were aggregated to create a monthly panel dataset spanning from July 2020 to July 2022.
@@
Our study leverages the natural experiment created by the launch of GitHub Copilot as a technical preview. Specifically, we exploit Copilot’s early-stage language endorsement, which included Python, JavaScript, Ruby, TypeScript, and Go, while excluding other comparable languages such as R, C, C#, C++, Java, PHP, and Scala. We select these non-Copilot-endorsed languages as the control group because they offer comparable functionality and are among the most frequently used languages in Microsoft-owned repositories. Our identification strategy, which contrasts programming languages across treatment and control groups, is consistent with prior studies (Yeverechyahu et al. 2024).
@@
Anecdotal evidence [^12] https://www.mckinsey.com/industries/technology-media-and-telecommunications/our-insights/thomas-dohmke-on-improving-engineering-experience-using-generative-ai and our interviews with Microsoft employees indicates that during the technical preview, access to Copilot was restricted to selected GitHub contributors, specifically employees of Microsoft/GitHub and maintainers of popular repositories who volunteered to participate in the restricted preview of the tool. This restricted access suggests that the primary users of Copilot during the technical preview were likely Microsoft and GitHub (organization) employees, ensuring that the tool was tested internally before its broader release. [^13] Dog Fooding; is Microsoft speak for internal use of their own software to ensure it is tested before public release https://devblogs.microsoft.com/oldnewthing/20110802-00/?p=10003
@@
Since Copilot usage cannot be identified at the individual level, we focus on Microsoft-owned repositories and contributors who actively contributed to these repositories during the observation period. By doing so, we aim to capture changes in OSS contributor behavior driven by Copilot, as contributors to Microsoft-owned repositories are more likely to have access to the tool during the technical preview period.
3.2 Data Collection
@@
We use GitHub’s API service to collect all the data for this research, enabling efficient and precise querying of repository / individual activities. This approach enables us to gather detailed information on measurements such as PRs, commits, reviews, authors, and repository / individual metadata, providing a comprehensive dataset for analyzing the impact of AI-assisted coding on OSS project development and maintenance activities.
At the project level, we define the treatment group as repositories whose primary programming language was among those endorsed by GitHub Copilot during the technical preview, while repositories using non-endorsed languages (e.g., C, C#, C++, Java, PHP, R, Scala) serve as the control group. Our project-level dataset consists of 2,755 repositories, with 1,660 in the Copilot-endorsed treatment group and 1,095 in the non-endorsed control group.
Our individual level dataset consisting of 1,699 contributors from GitHub, with 1,186 in the Copilot-supported treatment group and 513 in the non-Copilot-supported control group. [^14] There are 37,334 contributors who made at least one contribution to the repositories in our sample. The contributions include, for example, posting a comment, submitting a commit, or conducting a PR review. Among them, we filtered out 5,308 contributors who participated in more than three of the repositories we studied (we selected three repos to ensure sufficient variation for our PR reviewed repositories measure) . Then, we applied a programming language filter to construct a comparable treatment and control group for the individual-level analysis, and the data set was reduced to 1,699 contributors (who were users of the treated and control programming languages).
3.3 Variables
Table 2 presents the main variables and provide descriptive statistics. Our dataset is aggregated to the repository-month / contributor-month level, summary statistics are calculated based on repository-month / contributor-month observations.
Table 2: Variable Descriptions
Variable Description
Dependent Variables
Code Added (log) Line of codes submitted to a repository in a given month.
Commits (log) Commits submitted to a repository (by a contributor) in a given month. On GitHub, each commit typically represents a completed piece of work, such as fixing a bug, adding a feature, or improving existing code.
PRs (log) Pull requests submitted to a repository (by a contributor) in a given month. Each pull request represents a bundle of proposed changes (commits) that must be reviewed and approved before becoming part of the repository.
Techinal Debt Measured by PR rework (log), count of follow-up commits added to a PR after initial submission, capturing the extent of required code revision.
Maintenance Efforts Measured by PR review (log), count of PR reviews conducted by a contributor in a given month, analyzed at individual level.
Independent Variables
Treatment Indicator variable equal to 1 if the repository’s primary programming language is one of the five Copilot-supported languages, and 0 otherwise.
Copilot Indicator variable equal to 1 for months after the introduction of GitHub Copilot (July 2021), and 0 otherwise.
Treatment $\times$ Copilot Interaction term capturing the DiD estimate of Copilot’s effect on repository-level development and maintenance outcomes.
Control Variables
PRs (log) Included as a control in PR rework regressions to account for heterogeneity in code submission volume across repositories.
Month-year fixed effects Indicator variables capturing temporal shocks common across all repositories (e.g., seasonal patterns).
Repository fixed effects Repository-level constants controlling for time-invariant heterogeneity (e.g., project age, governance structure).
  • Note: All DVs are log-transformed. A value of one is added prior to transformation to address zeros.
3.3.1 Dependent Variable: PR Rework
In this study, we operationalize technical debt through process and outcome indicators, using rework rates, and maintenance effort, thereby connecting debt to developer behavior and organizational performance (Banker et al. 2021, Ramasubbu and Kemerer 2021). Specifically, Banker et al. (2021), Ramasubbu and Kemerer (2021) measure technical debt as the proportion of customized code in customer relationship management (CRM) systems that does not adhere to vendor-prescribed standards. Analogously, in the OSS context, we measure technical debt by the number of commit amendments (PR rework) required for a PR to satisfy repository standards, capturing the additional rework effort induced by nonconforming code.
On GitHub, a PR is a vehicle of contribution through which contributors participate in the development process. A PR typically contains one or more commits and often represents a "patch" or feature addition to the project. Any individual can submit a PR (a request to merge their contribution into the project), which is then (peer) reviewed by the core contributors, who have write access to the source code. The core contributors reviewing the PR can decide to merge the PR into the main project, request modifications or, reject it. If modifications are requested for a PR, the author of the PR can address the comments and modification requests and re-submit the code in the form of follow-up commits for another round of review. This revision process continues until all issues with the code are resolved and the code can be merged, or until the idea behind the code is no longer in alignment with the project goals and the PR is rejected. Taking this into account, we can measure the extent of rework done on a PR submitted by a contributor by identifying the number of commits that are added to the PR after its initial submission. We use the PR rework measure as a proxy to determine the technical debt attributable to initial code contributions submitted to the project and the maintenance-related efforts associated with the code contributions.
3.3.2 Dependent Variable: PR Review and PR Reviewed Repos
As described in the previous section, we use PR rework as a proxy for the accumulation of technical debt. Consistent with prior literature (Banker et al. 2021, Ramasubbu and Kemerer 2021), we further conceptualize PR reviews and PR reviewed repositories as measures of maintenance effort aimed at preventing or mitigating technical debt.
PR reviews capture the intensity of review activity undertaken by a contributor and are defined as the number of PR for which the contributor provides formal review during the observation period. This measure reflects the effort devoted to evaluating, debugging, and improving submitted code prior to integration.
PR reviewed repositories capture the breadth of a contributor’s maintenance responsibilities and are defined as the number of distinct repositories in which the contributor conducts at least one PR review during the observation period. This variable reflects the extent to which maintenance effort is distributed across multiple projects and repositories.
3.3.3 Moderation Variable: Core Contributors
For the conceptualization of core contributors we used contributors’ level of activity during the pretreatment period, measured by the number of commits, to classify contributors into core contributors (top 25%) and peripheral contributors (rest 75%) (Setia et al. 2012, Crowston et al. 2006). The classification of contributors allows us to examine whether the effect of Copilot varies depending on a core versus a periphery contributor. From Figure 3, we observe that core contributors perform the majority of development activities (both in terms of commits and PR) in the projects.
3.3.4 Descriptive Statistics
We report summary statistics for the key variables used in our analyses. Table 3 presents the descriptive statistics for the project-level analysis, including lines of code added, and the number of commits, PRs, and PR rework submitted; for the individual-level analysis, including the number of commits, PRs, PR reviews and PR reviewed repositories.
@@ 3.3.4 Descriptive Statistics
3.4 Estimation Models
3.4.1 Main Statistical Analysis
To estimate the effect of GitHub Copilot on repository / individual-level development and maintenance outcomes, we employed a DiD regression framework. This approach exploits the staggered introduction of Copilot across programming languages, comparing changes in outcomes for repositories / individuals use Copilot-supported languages (treatment group) with those use non-supported languages (control group) before and after Copilot’s technical preview. By examining within-repository / individual changes over time and contrasting them with contemporaneous changes in the control group, the DiD design isolates the impact of Copilot adoption from time-invariant repository characteristics and common temporal shocks.
The project level DiD model (repo i month t) is provided below:
@@ 3.4.1 Main Statistical Analysis
where $y_{i,t}$ refers to the outcome measures (development and maintenance) for project / individual $i$ in month $t$. $Copilot_{i,t}$ is the independent variable and a binary indicator that turns to 1 when the Copilot is released and functioned as our treatment. $CoreContributor_{i}$ is the moderator and a binary indicator that turns to 1 when one individual is identified as core contributor by pretreatment code contribution behaviour. The project-level / individual-level fixed effects are represented by $\gamma_{i}$, and $\delta_{t}$ represents the monthly fixed effects. $\epsilon_{i,t}$ is the robust standard error clustered at the project / individual level to account for the potential heteroskedasticity of the errors. The statistical results are presented in the appendix (Section 8.1).
3.4.2 Parallel Trends:
We followed prior research (e.g.,Autor 2003) to conduct an event study and test whether PR rework before the treatment are similar and parallel, an important assumption for the DiD framework. Specifically, the equations to test the parallel trends are:
$\text{PR rework}_{it}=\alpha+\sum_{\tau\neq-1}\beta_{\tau}\left(Copilot_{i}\times\mathds{1}\{t-T_{i}=\tau\}\right)+\gamma_{i}+\delta_{t}+\epsilon_{i,t}$ (3)
where ${PRrework}_{it}$ denotes the log number of PR rework actions in repository $i$ at time $t$; $Copilot_{i}$ indicates repositories associated with Copilot-supported languages; and $T_{i}$ denotes the Copilot introduction month. The coefficients $\beta_{\tau}$ capture the relative changes in rework activity $\tau$ months before or after treatment, using $\tau=-1$ as the omitted baseline. Repository fixed effects ($\gamma_{i}$) absorb time-invariant heterogeneity, while month-year fixed effects ($\delta_{i}$) account for aggregate shocks.
This event-study design enables us to visually and statistically evaluate the parallel-trends assumption and to characterize how Copilot’s effects unfold over time, offering a richer interpretation than a single aggregated DiD estimate.
4 Main Results
4.1 Project Level Analysis - Technical Debt
We analyzed the number of lines of code added, commits, and PRs as code development activities, and PR rework as a measurment of technical debt. Table 5 presents the results. After the introduction of GitHub Copilot, repositories whose primary programming language was endorsed by Copilot (treatment) experienced a significant increase in code development activities, such as the number of lines of code added increased 17.7% ($\exp(0.163)-1\approx 17.7\%$), commits increased 4.1% ($\exp(0.04)-1\approx 4.1\%$), and PRs increased 4.3% ($\exp(0.042)-1\approx 4.3\%$). This increase of productivity aligns with the analysis of industry experts (Peng et al. 2023) and the findings of recent academic research (Yeverechyahu et al. 2024).
Activity: Development Technical Debt
DV: Lines of Added Code Commits Pull Requests PR Rework
Copilot 0.163*** 0.04* 0.042*** 0.024**
(0.06) (0.022) (0.015) (0.01)
Project FE
Month FE
PR Controls
N 66,120 66,120 66,168 66,168
Adj. $R^{2}$ 0.516 0.603 0.691 0.81
More importantly, we find that the submitted PRs requires greater amount of rework, indicating a likely decrease in the quality of the submitted PRs. When analyzing the PRs submitted, the increased amount of code resubmissions is significant, even when controlling for the number of PRs, indicating an increased demand for code reviews per unit of code development. Specifically, we find that treatment repositories experienced 2.4% ($\exp(0.024)-1\approx 2.4\%$) more code rework, keeping the number of PRs submitted constant. These AI paired contributions are less likely to adhere to repository standards and therefore generate additional code rework effort, causing the accumulation of technical debt.
Figure 4 presents the event time analysis of the treatment effect on PR rework (leads-lags estimates). The coefficients for the pre-treatment periods are statistically insignificant, indicating that the parallel trends assumption holds (Angrist and Pischke 2009). The leads-lags estimate also indicate a long-term effect of Copilot. With the increasing deployment and adoption of AI, we expect more contributor engagement and frequent usage of Copilot. There is an upward trend in the post-treatment coefficients for PR rework that gradually becomes more positive over time (regression results are provided in the Supplementary Material). The observed pattern suggests a growing trend of increased code rework in repositories that supported AI coding partners.
Figure 4: Parallel trends and dynamic effects of the Copilot treatment for PR rework
4.2 Individual Level Analysis - Maintenance Efforts
We next examined commits and PRs as code development activities, and PR reviews (PR controlled) and PR reviewed repositories (PR controlled) as maintenance related activities. Based on our DiD analysis (Table 6), we find that core contributors engage in fewer development activities and in more maintenance activities after the launch of Copilot. The core contributors performed significant less code commits submissions, a decrease of 42.9% ($\exp(0.357)-1\approx 42.9\%$) compare to the rest of contributors. At the same time, the core contributors reviewed more PRs and do so across a broader set of repositories. This shift suggests that core contributors are not only stretching more of their limited time on maintenance tasks but also spreading themselves thinner across more repositories. The resulting burden of maintenance may come at the cost of reduced productivity among more experienced contributors.
Table 6: The impact of technical preview on GitHub’s core contributor behavior.
Activity: Development Maintenance
DV: Commits Pull Request PR Review PR Reviewed Repos
Copilot 0.142*** 0.091*** 0.04 0.008
(0.408) (0.03) (0.023) (0.012)
Core Contributor -0.357*** -0.027 0.06* 0.045**
$\times$ Copilot (0.052) (0.042) (0.035) (0.018)
Individual FE
Month FE
PR Controlled
N 51,408 51,408 51,408 51,408
Adj. $R^{2}$ 0.643 0.607 0.799 0.757
5 Robustness Check
We performed a series of tests to ensure that our findings are robust to various model specifications. Together, these tests increase confidence that our results are not driven by our model specification or unobserved variables.
@@ 5.1 Matching Results
As a robustness check, we employed Coarsened Exact Matching (CEM) to mitigate concerns about potential selection bias and ensure a more balanced comparison between treatment and control groups. CEM allows us to pre-process the data by matching units on a set of covariates that may influence both treatment assignment and outcomes. We matched repositories based on pretreatment code development characteristics, including Code Added, Commits, and PRs. After matching, the repository level dataset consisting of n = 2,510 repositories, with 1,486 in the treatment group and 1,024 in the control group. The treatment and control groups were more closely aligned in their baseline characteristics, reducing imbalance across key variables. We then re-estimated the DiD models using the matched sample. The results are listed in the table below. The results remain consistent with our main findings, reinforcing the conclusion that the integration of GitHub Copilot is associated with increased rework and maintenance related activities.
@@
Concept: Development Technical Debt